Circular No. 7180

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 K6 (LINEAR) Another apparently asteroidal object discovered by LINEAR (discov-

ery observation below; additional astrometry on MPEC 1999-K35), and posted on The NEO Confirmation Page, has been reported as cometary on May 24 by L. Šarounová (Ondřejov; faint coma with condensed nucleus) and by M. Tichý and Z. Moravec (Kleť; diffuse with coma diameter ≥ 10′′).

1999 UT  α_{2000}  δ_{2000}  m_{2}
May 20.22732  19^{h}01^{m}39.^{s}50  +9^{o}26′33′′.8  17.6

Preliminary parabolic orbital elements by B. G. Marsden:

T = 1999 July 24.130 TT  ω = 56.^{o}518
Ω = 245.426 }2000.0
q = 2.25099 AU  i = 46.403
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 22 19^{h}01.^{m}68 +10^{o}23.^{′}8 1.639 2.364 124.^{o}5 20.^{o}7 15.8
June 1 19 00.34 +15 48.7 1.567 2.332 127.9 20.1 15.7
11 18 56.47 +21 04.8 1.522 2.304 129.4 19.9 15.5
21 18 50.39 +25 54.0 1.502 2.283 129.0 20.2 15.5

V1333 AQUILAE C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence; and

P. Leisy and F. Patat, European Southern Observatory (ESO), report: “We have taken V -band and Cousins I-band CCD frames of V1333 Aql (= Aquila X-1) with the ESO 3.6-m telescope (+ EFOSC-2) in quiescence (Apr. 9) and near maximum brightness during the current outburst (May 21). The seeing (in I) was 0′′.8 and 1′′.1 in April and May, respectively, with a projected pixel size of 0′′.16. The barycenter of the image, corresponding to the variable object, shifts by 0′′.42 ± 0′′.03 to the west during activity. A similar effect is also seen in data taken with the Haute-Provence 1.2m telescope for the long-term monitoring program (pixel size 0′′.69 and seeing 1′′.6). This shows that the true optical counterpart of Aql X-1 is the interloper reported on IAUC 7086 to be located 0′′.46 to the west of the main object. This object has V = 21.6 and V − I_{c} = +2.2 (Chevalier et al. 1999, A.Ap., submitted). Combined analysis of this photometry and of low-resolution spectra taken previously with the same telescope suggest a late-K spectral type reddened by E(B − V) = 0.5 mag.”

1999 May 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green